OpenStudy (anonymous):

is 6 1/8 divided by 7 1/4 = 49/58 ?

OpenStudy (anonymous):

6 1/8 divided by 7 1/4 = 6 1/8 x 7 4/1= 42

OpenStudy (anonymous):

So no 42 is is the answer :)

OpenStudy (anonymous):

Do you get it?

OpenStudy (anonymous):

49/8 * 4/29 = 49/2 * 1/29 = 49/58 yes